ITES2001 Fourth Italian-Spanish conference on General Topology and its applications Bressanone, 27-30 June 2001

On abstract set-convergences

Alois Lechicki Fürth, Germany

This lecture deals with some abstract concepts of set-convergences. One of our principal interests is to establish connections between these convergences and classical ones, e.g. Vietoris or Hausdorff convergence.

2000 Mathematics Subject Classification. 54B20, 54A20.

Key words and phrases. Set-convergences, Vietoris convergence, Hausdorff convergence.